A flag pole 24 foot high is 10 foot away from the base of its anchor cable. How long is the anchor cable?
Answer by sharon1948(6)

Problem #287277
Pythagorean Theorem
A flag pole 24 foot high is 10 foot away from the base of its anchor cable. How long is the anchor cable?
I would draw out the triangle and label it for better understanding.
Let:
a=24
b=10
c=c^2
Formula:
a^2+b^2=c^2
Solve:
With the information you have rewrite the formula
24^2+10^2=c^2
Next find the exponential values of 24^2 and 10^2
(24 times 24 and 10 times 10)
576+100=c^2
Then find the sum of
576+100=676
Now find the square root of the sum 676.
c= 26
The length of the anchor cable is 26ft.